OpenStudy (anonymous):

if g(x)= 8-7x and h(x)= 6x + x squared, find g(-9)

OpenStudy (anonymous):

will give medal!

OpenStudy (anonymous):

simple put x=-9 in g(x)= 8-7x and take teh result

OpenStudy (anonymous):

g(-9)=8-7(-9) =8+72 =80